how long will it take a given sum of money to increase 4 times its present value when compounded half yearly at 7% rate of interest
1
Expert's answer
2020-10-05T11:00:38-0400

solution


let the present value "P=1" . Then the future value will be "A=4"

Since interest is compounded half yearly, "i=\frac{7\%}{2}=0.035"



"A=P*(1+i)^n""4=(1.035)^n""log (4)=n\ log (1.035)"

"n=\frac{log(4)}{log(1.035)} = 40.2976"

Since payments are made half yearly, the period it takes to grow an amount by a factor of 4 is


"=\frac{40.2976}{2}= 20.1488\ years"

answer: it will take 20.15 years